OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

The form is (x - h)^2 + (y - k)^2 = r^2

OpenStudy (anonymous):

I don't get it, switch would it be?

OpenStudy (anonymous):

Basically, r stand for radius. So if the radius is 8, where would gthe radius go and what would happen to it based on the form?

OpenStudy (anonymous):

Then switch the signs of the center and place them where h and k are. h with the x and k with the y

OpenStudy (anonymous):

So would it be (x – 2)2 + (y + 1)2 = 8?

OpenStudy (anonymous):

Remember, it's r^2 not r. It would be (x – 2)^2 + (y + 1)^2 = 64